If a recessive condition occurs in 16% of individuals in a population, what will be the frequency of occurrence of the condition in the next generation assuming the affected individuals all die before reproducing?
Answer #1

Since the recessive condition affects 16% of the current generation,

Then Frequency(Recessive Homozygotes) = q2 = 0.16

=> q = 0.4
=> p = 1-q = 0.6
=> 2pq = 2 * 0.6 * 0.4 = 0.48

Thus the current generation is comprised of 16% Homozygous recessive individuals (who do not reproduce), 48% heterozygotes and 36% Homozygous Dominant individuals.

Now, sine the affected individuals die without reproducing, the net generation can have affected individuals only by the mating of heteroygotes, both of whom pass on the recessive allele 1/2 of the time. In other words, 1/4th of all progeny borne of Heterozygotes in the present generation will be Homozygous recessive for the condition.

Now, in the remaining population(Comprised of Homozygous dominant and Heterozygous individuals), 0.48/(0.48+0.36) = 0.84 are Heterozygotes and 0.16 are Homozygous dominant.

Thus, frequency of affected individuals (homozygous recessive) in next generation;

= Fraction of Progeny produced by two Heterozygotes mating * 1/4
= 0.84 * 0.84 * 0.25
= 0.1764

Thus, 17.64 % of individuals in the next generation will be affected.